Write the chemical reaction for the autoionization of water (equilibrium constant Kw). Look up the numerical...

  1. Write the chemical reaction for the autoionization of water (equilibrium constant Kw).

  2. Look up the numerical value for Kw at Tinitial (Use the average of Tinitial from both trials) -Its 6.81x10^-15

  3. How is the equation in Question 1 related to the net ionic equation for the strong acid-strong base neutralization shown below?

    H3O+(aq) + OH-(aq) → H2O(l) + H2O(l)

  4. Use the value of Kw and the relationship between the equations in Questions 1 and 3 to find the numerical value

    of K for the neutralization reaction in question 3.

  5. Use the value of K calculated for the neutralization reaction in Question 4 to calculate ΔG for the neutralization reaction. (Use average Tinitial from the two trials as the value of T.)

  6. Use the values of ΔG and ΔHn that you have already calculated to calculate ΔS° for the reaction. (Use average Tinitial from the two trials as the value of T.)
Answer #1

Chemical reaction for autoionization of water, H₂O(l) + H₂O(l) H₂O + (aq) + OH (94) Kw = 6.81810-15 = [H₂0 tags] [ot cons] Ne